# Hybrid Quantum Physics-informed Neural Network: Towards Efficient Learning of High-speed Flows

### v1

@@ 4 Conclusion
Here we discuss the implementation of a HQPINN model for high-speed flow problems, compared to classical and quantum PINN models. A qualitative summary is provided in Table 4.
For problems with harmonic solutions (Sec. 3.1), quantum models achieve high solution accuracies using very few trainable parameters. Classical models trains faster but are less economical in parameter count. With sufficient parameterization, hybrid models may outperform both quantum and classical models (see Fig. 3d).
For problems with non-harmonic solutions (Sec. 3.2), classical models significantly outperform quantum models. Hybrid models correct for overshoot and under-damped oscillations presented by quantum models (Fig. 5), but are only intermediate in terms of accuracy as they suffer from large variance in training loss due to classical and quantum sub-models. With sufficient parameterization, hybrid models may be able to match the performance of classical models (Fig. 6d). Trainability varies from medium–high (Sec. 3.2) to low (Sec. 3.3), depending on the complexity of the problem.
In conclusion, the HQPINN excels at problems with harmonic solutions, combining the best accuracy, trainability and scalability features from both classical and quantum networks. However, the HQPINN appears ill-suited for difficult problems with shocks and discontinuous solutions, with poor trainability for more complex problems. Hybrid PINN architectures may find use in applications where the nature of the solution is not known a-priori. In practice, a hybrid classical-quantum architecture can be implemented in a CPU-QPU environment, leveraging the capabilities of the latest quantum chips and processors in the pipeline, such as Google’s Willow [Acharya2024] and Microsoft’s Majorana I [Aghaee2025]. Specifically, the quantum processing unit (QPU) can efficiently process the quantum layer for harmonic components, while the CPU handles the classical layer for non-harmonic components. Future work may include adaptive weight balancing between classical and quantum sub-network layers, use of GPU-accelerated classical simulation of quantum models and fault-tolerant quantum hardware as they become available.
Model Harmonic Non-harmonic Number of trainable parameters
Accuracy Trainability Accuracy Trainability
Classical Medium High High High High
Quantum High Low Low Low Low
Hybrid Medium–High Medium–High Medium Varies Medium–High
Table 4: Qualitative assessment of PINN models in the present study.
@@ Acknowledgements
This research is supported by the National Research Foundation, Singapore and the Agency for Science, Technology and Research (A*STAR) under its Quantum Engineering Programme (NRF2021-QEP2-02-P03), and A*STAR C230917003.
@@ A.1 Parameterized quantum circuit design
The performance of QML depends on the expressivity and trainability of the model [Abbas2021], and even within the simple design framework of alternating feature map and ansatz layers, there are many configurations to choose from.
Following previous works[Kordzanganeh2023, xiao2024physics], we employ a four-qubit PQC as an entangled $RZ-RX-RZ$ gates for an ansatz layer $A(\theta)$, and a $RY$ gates for a feature map layer $S(\varphi)$ (Fig. A1). In this study, we consider a PQC with the following gate sequence: $A-(S-A)_{1}-\dots-(S-A)_{l-1}$.
[…]
@@ A.2 Damped harmonic oscillator
The quantum Fourier model (Eq. 9) could extract harmonic features from a suitable dataset [Kordzanganeh2023] and solve PDEs with harmonic features [xiao2024physics]. Here, we test the use of a hybrid classical-quantum model on a simple problem of a 1D damped harmonic oscillator, which is also an introductory PINN problem [PINN_intro]. The displacement $u$ in time $t$ is described by a 1D second-order ordinary differential equation (ODE) as,
@@
Following [PINN_intro], we set $m=1$, $\mu=4$, $k=400$, $\lambda_{1}=10^{-1}$ and $\lambda_{2}=10^{-4}$ and implement PINN on Pytorch [Paszke2019] and Pennylane [Bergholm2018] using Adam optimizer with initial learning rate 0.002 up to $2000$ epochs. The parallel neural network layers in this test are combinations of the following:
@@
As proof of concept, we tested three parallel HQPINN configurations, namely classical–classical, quantum–quantum and hybrid quantum–classical neural networks [Kordzanganeh2023]. Figure A2 shows that classical–classical networks struggles to learn the damped oscillations and requires more iterations to achieve convergence ($\approx 15,000$ epochs [PINN_intro]). The quantum–quantum network converges with high fidelity, demonstrating the power of the quantum Fourier model in modeling harmonic solutions [schuld2021effect, xiao2024physics]. Notably however, the quantum model is worse than classical model at early data fitting. By combining complementary features from classical and quantum layers, the hybrid model outperforms classical and quantum neural networks (Fig. A2c).
@@ A.3 Computational Fluid Dynamics simulation
Computational Fluid Dynamics (CFD) simulation of steady-state transonic flow (Eq. 1 and 2) past a NACA2012 airfoil is implemented on Ansys Fluent 2022 R2 [ansys_fluent] using pressure-based solver on inviscid pressure-velocity coupled scheme. For spatial discretization, we apply the least squares cell-based discretization for gradients, second-order discretization for pressure, and second-order upwind discretization for density, momentum, and energy. Residual of $10^{-6}$ is set as the convergence criterion for all variables.
Figure A3 shows the computational domain for a 2D compressible flow over a NACA0012 airfoil with 1 m span in a 4.5 m by 4.5 m square grid. Periodic boundary conditions are applied to both the top and bottom of the domain. We specify an inlet airflow velocity at Mach number 0.8 and a temperature 288.15 K. The outlet boundary is set at null air pressure and a temperature of 288.15 K. Free slip condition is applied at the surface of the airfoil. The finite volume grid consists of 25 000 cells with mesh refinement so that cell widths are approximately 30 mm in the domain and 10 mm near the airfoil surface.
Figure A3: (Left) Computational domain for modeling a transonic flow past a NACA0012 airfoil with 1 m span in a 4.5 m by 4.5 m square grid. (Right) CFD solutions show transonic shock regions for (a) Mach number, (b) temperature, (c) density and (d) pressure.
Figure A3(a–d) illustrates the distribution of airflow properties, including Mach number, temperature, density, and pressure. The far field subsonic flow at Mach 0.8 transitions to transonic behavior as it passes over the NACA0012 airfoil, generating a normal shock near the trailing edge. Away from the stagnation region, the pressure becomes negative following momentum conservation. Both air density and temperature distributions are similar to air pressure, since they are directly proportional to pressure based on the ideal gas law.

### v2

@@ 3.2 Discontinuous Euler Equation (Contact Discontinuity)
3.4 Summary of Results
This section highlights key insights based on model performances across the three benchmark problems. Here, we benchmark trainability using the epoch ratio defined as the ratio of number of epochs required to reach a specified threshold loss and total number of simulated epochs; the threshold loss are set as {Harmonic: $10^{-5}$, Non-harmonic: $10^{-4}$, Transonic: 101010} for the respective problems. Also, we benchmark costs using the parameter ratio defined as the ratio of number of parameters of the model and that of the classical model of the problem. Table 1 summarizes loss, epoch ratio, and parameter ratio for representative configurations of classical, quantum, and hybrid PINNs.
Harmonic (Sec. 3.1) Size Loss Epoch ratio Parameter ratio
Classical 10-4 $\sim 10^{-7}$ 0.260 1.000
Quantum 2 $\sim 10^{-8}$ 0.675 0.146
Hybrid 10-4-2 $\sim 10^{-7}$ 0.350 0.573
Non-harmonic (Sec. 3.2) Size Loss Epoch ratio Parameter ratio
Classical 20-4 $\sim 10^{-7}$ 0.310 1.000
Quantum 2 $\sim 10^{-4}$ $>$ 1 0.056
Hybrid 20-4-2 $\sim 10^{-6}$ 0.380 0.522
Transonic (Sec. 3.3) Size Loss Epoch ratio Parameter ratio
Classical 40-4 $\sim 10^{-2}$ 0.310 1.000
[…]
Hybrid 40-4-2 $\sim 10^{-1}$ 0.725 0.509
Table 1: Benchmarked performance of classical, quantum and hybrid PINN models in terms of loss, epoch ratio and parameter ratios (definitions in text), based on representive model size for each architecture.
Key findings:
  1. 1.
    Quantum PINNs excel in harmonic regimes, offering high accuracy with very low parameter costs (Table 1) due to the Fourier structure of PQCs (Section 3.1). However, quantum models have low trainability and struggle to generalize to discontinuities or shocks (Sections 3.23.3).
  2. 2.
    HQPINNs can mitigate spurious artifacts, including overshoot near discontinuities (Fig. 5), by leveraging classical sub-networks to handle non-harmonic features.
  3. 3.
    HQPINNs can match the best performance of classical models even for non-harmonic problems (Fig. 6d), at reduced parameter costs. This suggests that, in certain cases, the quantum sub-network could complement the training of models with non-harmonic features.
  4. 4.
    HQPINNs are highly trainable for both harmonic and non-harmonic problems as shown by the respective epoch ratios (Table 1). Note however that this trainability is reduced if the quantum layer were under-parameterized (Fig. 8).
  5. 5.
    Quantum emulation cost is a constraint. Simulating PQCs classically results in per-epoch runtimes up to 100× slower than fully classical models (Table A3, Appendix A.5); scalability is a practical concern on current hardware.
Overall, HQPINNs can potentially offer competitive performance at economical parameter costs across both harmonic and non-harmonic problem types, without falling decisively short in any single performance metric (Table 1).
@@ 4 Conclusion
Here we briefly review and discuss the implementation of a HQPINN model for high-speed flow problems, compared to classical and quantum PINN models.
For problems with harmonic solutions (Sec. 3.1), quantum models easily achieve high solution accuracies using very few trainable parameters but with low trainability. For problems with non-harmonic solutions (Sec. 3.2), classical models are significantly more accurate and trainable than quantum models but this also comes at a significant parameter cost. Either way, hybrid models are balanced in all aspects of accuracy, trainability and efficiency (Table 1).
In conclusion, the HQPINN is a highly dependable, problem agnostic architecture that can adapt useful features from both classical and quantum sub-networks. Based on our benchmarked study, we find that the HQPINN is currently limited by the relatively small scale of quantum sub-network which is unable to cope with more complex shock features of the aerofoil problem, which degrades the overall performance. Nevertheless, with rapid progress in quantum technologies, larger scale quantum hardware will eventually be realized and we believe hybrid PINN architectures will eventually find use in broad applications, especially where the nature of the solution is not known a-priori.
In practice, a hybrid classical-quantum architecture can be implemented in a CPU-QPU environment, leveraging the capabilities of the latest quantum chips and processors in the pipeline, such as Google’s Willow [30] and Microsoft’s Majorana I [31]. Realistically, the quantum processing unit (QPU) can efficiently process the quantum layer for harmonic components, while the CPU handles the classical layer for non-harmonic components. Future work may include adaptive weight balancing between classical and quantum sub-network layers, use of GPU-accelerated classical simulation of quantum models and fault-tolerant quantum hardware as they become available.
@@ Acknowledgements
This research is supported by the National Research Foundation, Singapore and the Agency for Science, Technology and Research (A*STAR) under Quantum Engineering Programme (NRF2021-QEP2-02-P03), Hybrid Quantum-Classical Computing HQCC 1.0 (SC23/24-7534CI) and A*STAR C230917003.
@@ A.1 Parameterized quantum circuit design
The performance of QML depends on the expressivity and trainability of the model [9], and even within the simple design framework of alternating feature map and ansatz layers, there are many configurations to choose from [17].
Following previous works[13, 16, 17], we employ a four-qubit PQC as an entangled $RZ-RX-RZ$ gates for an ansatz layer $A(\theta)$, and a $RY$ gates for a feature map layer $S(\varphi)$ (Fig. A1). In this study, we consider a PQC with the following gate sequence: $A-(S-A)_{1}-\dots-(S-A)_{l-1}$.
@@ A.2 Damped harmonic oscillator
The quantum Fourier model (Eq. 9) could extract harmonic features from a suitable dataset [13] and solve PDEs with harmonic features [16, 17]. Here, we test the use of a hybrid classical-quantum model on a simple problem of a 1D damped harmonic oscillator, which is also an introductory PINN problem [32]. The displacement $u$ in time $t$ is described by a 1D second-order ordinary differential equation (ODE) as,
@@
Following [32], we set $m=1$, $\mu=4$, $k=400$, $\lambda_{1}=10^{-1}$ and $\lambda_{2}=10^{-4}$ and implement PINN on Pytorch [24] and Pennylane [25] using Adam optimizer with initial learning rate 0.002 up to 200020002000 epochs. The parallel neural network layers in this test are combinations of the following:
@@
As proof of concept, we tested three parallel HQPINN configurations, namely classical–classical, quantum–quantum and hybrid quantum–classical neural networks [13]. Figure A2 shows that classical–classical networks struggles to learn the damped oscillations and requires more iterations to achieve convergence (15,000$\approx$ 15,000≈ 15 , 000 epochs [32]). The quantum–quantum network converges with high fidelity, demonstrating the power of the quantum Fourier model in modeling harmonic solutions [10, 16, 17]. Notably however, the quantum model is worse than classical model at early data fitting. By combining complementary features from classical and quantum layers, the hybrid model outperforms classical and quantum neural networks (Fig. A2c).
@@ A.3 Computational Fluid Dynamics simulation
Computational Fluid Dynamics (CFD) simulation of steady-state transonic flow (Eq. 1 and 2) past a NACA2012 airfoil is implemented on Ansys Fluent 2022 R2 [33] using pressure-based solver on inviscid pressure-velocity coupled scheme. The inviscid flow assumption retains key features such as shock waves, and is commonly employed in PINN studies [3, 29].
For spatial discretization, we apply the least squares cell-based discretization for gradients, second-order discretization for pressure, and second-order upwind discretization for density, momentum, and energy. Residual of $10^{-6}$ is set as the convergence criterion for all variables.
Figure A3 shows the computational domain for a 2D compressible flow over a NACA0012 airfoil with 1 m span in a 4.5 m by 4.5 m square grid. Periodic boundary conditions are applied to both the top and bottom of the domain. We specify an inlet airflow velocity at Mach number 0.8 and a temperature 288.15 K. The outlet boundary is set at null air pressure and a temperature of 288.15 K. Free slip condition is applied at the surface of the airfoil. The finite volume grid consists of 25 000 cells with mesh refinement so that cell widths are approximately 30 mm in the domain and 10 mm near the airfoil surface. Mesh convergence study is conducted.
Figure A3: (a) Computational domain for modeling a transonic flow past a NACA0012 airfoil with 1 m span in a 4.5 m by 4.5 m square grid. Compare (b) normalized mean pressure field over airfoil and (c) normalized pressure along surface chord, between current simulation and reference [34].
Figure A3(b) shows normalized mean pressure field over airfoil for current simulation and the reference [34]. The far-field subsonic flow at Mach 0.8 transitions to transonic behavior as it passes over the NACA0012 airfoil, creating a normal shock near the trailing edge. Figure A3(c) compares normalized pressure spanning the airfoil chord. Our results from the current simulation exhibit good agreement with reported values [34].
@@ A.4 Quantum and hybrid models for transonic flow
A.5 Tables for performance benchmarking
Model Size Trainable parameters Loss Density error Pressure error
cc 10–4 845 $6.13\times 10^{-7}$ $1.59\times 10^{-6}$ $3.28\times 10^{-9}$
10–7 1505 $\mathbf{3.16\times 10^{-7}}$ $7.33\times 10^{-7}$ $2.22\times 10^{-9}$
20–4 2845 $6.65\times 10^{-7}$ $2.02\times 10^{-6}$ $3.96\times 10^{-9}$
qq 2 87(+36) $\mathbf{2.63\times 10^{-8}}$ $9.76\times 10^{-9}$ $1.01\times 10^{-9}$
3 87(+54) $1.71\times 10^{-7}$ $1.34\times 10^{-7}$ $1.87\times 10^{-9}$
4 87(+72) $2.75\times 10^{-7}$ $2.80\times 10^{-7}$ $2.46\times 10^{-9}$
hy 10–4–2 448(+36) $\mathbf{1.98\times 10^{-7}}$ $2.29\times 10^{-7}$ $8.44\times 10^{-9}$
10–7–2 778(+36) $3.31\times 10^{-7}$ $3.24\times 10^{-7}$ $9.04\times 10^{-9}$
20–4–2 1448(+36) $8.62\times 10^{-7}$ $3.19\times 10^{-7}$ $4.98\times 10^{-9}$
Table A1: Smooth Euler Equation (Section 3.1). Trainable parameter count excludes quantum parameter count in parenthesis. Mean loss and relative error for density and pressure at the 20 000th epoch, averaged over 6 runs. Best loss results are shown in bold.
Model Size Trainable parameters Loss Density error Pressure error
cc 10–4 845 $\mathbf{6.89\times 10^{-7}}$ $2.89\times 10^{-4}$ $1.92\times 10^{-6}$
10–7 1505 $1.39\times 10^{-6}$ $4.07\times 10^{-4}$ $6.06\times 10^{-7}$
20–4 2845 $1.51\times 10^{-6}$ $4.67\times 10^{-4}$ $6.59\times 10^{-7}$
qq 2 87(+36) $5.00\times 10^{-4}$ $1.20\times 10^{-3}$ $4.51\times 10^{-5}$
3 87(+54) $1.45\times 10^{-4}$ $1.30\times 10^{-3}$ $1.53\times 10^{-5}$
4 87(+72) $\mathbf{1.39\times 10^{-4}}$ $9.65\times 10^{-4}$ $1.24\times 10^{-5}$
hy 10–4–2 448(+36) $7.53\times 10^{-6}$ $5.29\times 10^{-4}$ $1.96\times 10^{-6}$
10–7–2 778(+36) $3.93\times 10^{-5}$ $3.55\times 10^{-4}$ $3.52\times 10^{-5}$
20–4–2 1448(+36) $\mathbf{6.98\times 10^{-6}}$ $2.37\times 10^{-4}$ $3.29\times 10^{-6}$
Table A2: Discontinuous Euler Equation (Section 3.2). Trainable parameter count excludes quantum parameter count in parenthesis. Mean loss and relative error for density and pressure at the 20 000th epoch, averaged over 6 runs. Best loss results are shown in bold.
Model Size Trainable parameters Time (s/epoch) Loss Pressure error
cc 40–4 10 628 0.11 $1.11\times 10^{-2}$ $1.35\times 10^{-3}$
40–7 20 468 0.21 $\mathbf{2.48\times 10^{-3}}$ $6.79\times 10^{-4}$
80–4 40 388 0.20 $3.13\times 10^{-3}$ $7.74\times 10^{-4}$
qq 2 140(+48) 1.9 $9.67\times 10^{1}$ $3.59\times 10^{-3}$
4 140(+96) 4.6 $8.48\times 10^{1}$ $2.53\times 10^{-3}$
6 140(+144) 9.0 $\mathbf{6.65\times 10^{1}}$ $4.21\times 10^{-3}$
hy 40–4–2 5360(+48) 1.3 $1.36\times 10^{-1}$ $4.05\times 10^{-3}$
40–7–2 10 280(+48) 1.1 $3.71\times 10^{0}$ $1.95\times 10^{-3}$
80–4–2 20 240(+48) 1.1 $\mathbf{2.56\times 10^{-2}}$ $2.46\times 10^{-3}$
Table A3: 2D Transonic Aerofoil Flow (Section 3.3). Trainable parameter count excludes quantum parameter count in parenthesis. Average wall time shown in seconds per epoch. Mean loss and relative error for pressure averaged over 6 runs at the 42 000th epoch. Best loss results are shown in bold.